what percent of 1800 is 400?
Found 4 solutions by ichudov, chawki, amalm06, ikleyn:
Answer by ichudov(507)

there is a calculator for this on my site.

Answer by chawki(7)

1800 -> 100%
400 -> x
x = (400*100) / 1800
x = 22.22 %

Answer by amalm06(224)

(400/1800)*100=(2/9)(100)=200/9% or 22.2% (Answer)

Answer by ikleyn(53937)